OpenStudy (anonymous):

x=7 9/6 y =11 6 /12 xy=?

hero (hero):

First reduce the fractions: 9/6 = 3/2 6/12 = 1/2 Then multiply: (7 + 3/2)(11 + 1/2)

OpenStudy (anonymous):

ok thx :)

hero (hero):

Do you know how to multiply that?